Concentrated h2so4 has density of 1.9g/ml and is 99% by weight . Calculate molarity of h2so4

Grade:12

2 Answers

Yash Yash
4 years ago
Molar mass of h2so4-98gm
Density is 1.9g/ml
Which means 1000 ml(1lire) h2so4 will weigh 1900gm
% w/w-99% 
i.e. H2so4 in 1000 ml is 99/100*1900=1881gm
No. Of moles of h2so4 is1881÷19=19.19
Therefore, molarity=no. Of moles of solute/volume of soln. 
Molarity=19.19/1
=19.19
Khimraj
4 years ago
molarity = (given mass * density * 10)/(molar mass)
 = 99*1.9*10/98 = 19
Hope it clears......................................................